# PrivacyRestore: Privacy-Preserving Inference in Large Language Models via Privacy Removal and Restoration

### v3

@@ Abstract
The widespread usage of online Large Language Models (LLMs) inference services has raised significant privacy concerns about the potential exposure of private information in user inputs to malicious eavesdroppers. Existing privacy protection methods for LLMs suffer from either insufficient privacy protection, performance degradation, or large inference time overhead. To address these limitations, we propose PrivacyRestore, a plug-and-play method to protect the privacy of user inputs during LLM inference. The server first trains restoration vectors for each privacy span and then release to clients. Privacy span is defined as a contiguous sequence of tokens within a text that contain private information. The client then aggregate restoration vectors of all privacy spans in the input into a single meta restoration vector which is later sent to the server side along with the input without privacy spans. The private information is restored via activation steering during inference. Furthermore, we prove that PrivacyRestore inherently prevents the linear growth of the privacy budget. We create three datasets, covering medical and legal domains, to evaluate the effectiveness of privacy preserving methods. The experimental results show that PrivacyRestore effectively protects private information and maintain acceptable levels of performance and inference overhead.
@@
Ziqian Zengthanks: Equal contribution.  thanks: Corresponding author  1, Jianwei Wang11footnotemark: 1  1, Junyao Yang1, ZhengdongLu1, Huiping Zhuang1, Cen Chen1,2 1South China University of Technology 2 Pazhou Laboratory, China    zqzeng@scut.edu.cn,    wiwjwilliam@mail.scut.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) have emerged as powerful tools in various domains, including healthcare (Chen et al., 2023; Xu et al., 2023), law (Wu et al., ; Deng et al., 2023a), and finance (Wu et al., 2023; Xie et al., 2023). With the exception of a very small portion of users who have the resources and expertise to deploy LLMs locally, the vast majority of users access and interact with these powerful models through online inference services.
@@
However, the widespread usage of online LLMs inference services has raised significant privacy concerns, especially regarding the potential risk of private information being leaked through user inputs when interacting with LLMs deployed on cloud platforms. User inputs often contain sensitive information such as details in medical records and legal cases. Potential threats may arise from eavesdropper attackers intercepting user queries during transmission to cloud platforms for malicious purposes. For example, in sensitive domains like medical diagnosis, if a user’s input containing personal health information, such as "I was previously diagnosed with HIV, and lately I’ve been experiencing fever and diarrhea…" is disclosed, it may cause troubles to their life.
@@
In this paper, we focus on protecting the private information contained in user inputs during LLM inference stage. In this setting, the client submits inputs to the server (also known as the service provider) and there is a risk that inputs might be disclosed by attackers. Current methods for protecting user inputs can be categorized into two categories: Secure Multi-Party Computation (SMPC) and Differential Privacy (DP). SMPC based methods (Hao et al., 2022a; Li et al., 2023a; Liang et al., 2024) utilize encryption protocols and algorithms to enable collaborative computation without revealing original data to others. However, SMPC methods have large inference time overhead, making them impractical for real-time applications. For example, running a single pass inference on the RoBERTa-Base (Liu et al., 2019) requires 168.43 seconds (Hao et al., 2022b). DP based methods (Feyisetan et al., 2020, 2019; Xu et al., 2020; Bo et al., 2021) introduce the definition of $d_{\chi}$-privacy (Chatzikokolakis et al., 2013; Alvim et al., 2018) and apply a word-level text-to-text privatization on data locally before transmitting data to the server. Nevertheless, DP based methods inevitably degrade the performance of downstream tasks due to noise injection, which is known as the privacy-utility trade-off. Additionally, as text length increases, the performance degradation becomes pronounced, This phenomenon is known as the linear growth of the privacy budget Mattern et al. (2022b) in differential privacy. Hence, there is a need to develop privacy-preserving methods which can effectively safeguard the privacy of user inputs while maintaining high-quality outputs, without incurring prohibitive computational costs.
We propose PrivacyRestore which directly removes privacy spans in user inputs and restores private information via activation steering (Li et al., 2023c; Turner et al., 2023; Hernandez et al., 2023) during model inference. Our method is based on two key assumptions: (a) Private information is confined within specific a contiguous sequence of tokens, termed “privacy span”, rather than being dispersed throughout the entire input. Privacy span is defined as a contiguous sequence of tokens within a text that contain private information. The removal or proper redaction of privacy spans significantly impedes unauthorized parties from reconstructing or inferring the underlying private information. For instance, if privacy spans "HIV" , "fever" and "diarrhea" are removed from a medical record "I was previously diagnosed with HIV, and lately I’ve been experiencing fever and diarrhea…", attackers can not recover any private information. (b) In a particular domain, the number of potential privacy spans is limited and finite. For example, in the application of medical diagnosis, privacy spans generally pertain to symptoms and disease names, and the number of possible symptoms and disease names is inherently limited. Moreover, although medical knowledge and terminology inevitably evolve, the core set of symptoms and diseases remains relatively stable and finite.
@@
PrivacyRestore operates in two stages: the preparation stage and the inference stage. In the preparation stage, we first identifies the attention heads where the activation steering occurs. Second, each privacy span is encoded to a vector named restoration vector. This stage is conducted on the server side. Our method is plug-and-play, requiring only the restoration vectors to be trainable, while keeping the LLM frozen. Once training is complete, the users keep all restoration vectors on the client side. In the inference stage, the user construct a meta vector by first estimating the importance of each privacy span in the input and then calculating a weighted sum of the corresponding restoration vectors. The user then removes the privacy spans from the input and submits the remaining input along with the meta vector to the server. The server uses the meta restoration vector to restore the removed privacy spans through activation steering.
@@
To prevent the leakage of privacy spans via reverse-engineering on the meta vector, the $d_{\chi}$-privacy mechanism (Feyisetan et al., 2020) is applied on the meta vector before transmission on the client side. $d_{\chi}$-privacy mechanism is a variant of the differential privacy mechanism (Dwork et al., 2016). By applying $d_{\chi}$-privacy to the meta restoration vector instead of words, our method inherently addresses the issue of linear growth of privacy budget (Mattern et al., 2022a) commonly encountered in $d_{\chi}$-privacy and other DP variants. Experimental results demonstrate that the proposed method effectively protects private information and maintains satisfactory performance and inference efficiency.
The contributions of our paper are summarized as follows,
@@ 2 Related Works
2.1 Secure Multi-Party Computation (SMPC)
Secure multi-party computation (SMPC) methods utilize multi-party encryption algorithms to enable collaborative computation among multiple parties while protecting the privacy of their data. However, most nonlinear operations in LLMs cannot directly support secure multi-party computation. To address this challenge, current SMPC methods focus on two optimization directions: model structure-oriented optimization and protocol-oriented optimization.
The model structure-oriented approach aims to replace SMPC-unfriendly nonlinear operations with SMPC-friendly alternatives. For instance, MPC-Former (Li et al., 2023a) approximates nonlinear operations in Transformer using polynomials and maintains performance through model distillation. MERGER (Liang et al., 2024) integrates previous techniques to natural language generation (NLG) tasks by bypassing embedded computation and reorganizing linear operations in Transformer modules, further enhancing computational efficiency and model performance. In contrast, the protocol-oriented approach focuses on designing efficient SMPC operators for nonlinear operations in LLMs while preserving the original model structure. Recent works Hao et al. (2022a); Liu and Liu (2023); Zheng et al. (2023b); Gupta et al. (2023) have improved the efficiency of nonlinear operations in privacy-preserving LLMs inference by utilizing various SMPC protocols, such as confusion circuit and function secret sharing.
Although SMPC-based methods can be applied to protect user inputs during model inference, they still suffer from large inference time overhead. For example, inference on the RoBERTa-Base model takes 168.43 seconds (Hao et al., 2022b), making current SMPC methods impractical for online LLM inference services.
2.2 Differential Privacy (DP)
Differential Privacy (DP), as introduced by Dwork et al. (2016), is designed to protect individual privacy by preventing attackers from identifying specific participants in a dataset. Several variants of DP have been developed to enhance privacy protection across various settings, adapting the core principles of DP to different types of data and threat models. Notable examples include Centralized Differential Privacy (CDP), Local Differential Privacy (LDP) and $d_{\chi}$-privacy.
CDP (Dwork et al., 2016) operates under the assumption that all data has been stored in a central repository. It guarantees that attackers cannot distinguish between any two adjacent repositories based on query results. In contrast, LDP (Duchi et al., 2013) provides a stronger guarantee, ensuring that attackers cannot distinguish between any two adjacent inputs. Mattern et al. (2022b) and Utpala et al. (2023) propose using paraphrasing techniques to achieve LDP on user inputs. The formal definitions of CDP and LDP are provided in Appendix D.
LDP allocates the same privacy budget $\epsilon$ to all adjacent input pairs, regardless of their similarity. Applying the same $\epsilon$ forces each user input to be indistinguishable from any other, which can negatively impact data utility. However, it is sufficient for privacy protection to make each user input indistinguishable only from its closer counterparts. To address this, $d_{\chi}$-privacy (Feyisetan et al., 2019), a relaxed version of LDP, incorporates metrics that measure the similarity between inputs, allowing for more flexible control over the privacy budget. $d_{\chi}$-privacy is defined as,
Definition 2.1.
($\bm{d_{\chi}}$-privacy). A randomized mechanism $\mathcal{M}:\mathcal{I}\to\mathcal{O}$ fulfills $\epsilon$-$d_{\chi}$-privacy if for all adjacent inputs $I,I^{\prime}\in\mathcal{I}$ and all possible outputs $O\subset\mathcal{O}$,
$\quad\mathbb{P}\left(\mathcal{M}(I)\in O\right)\leq\exp(\epsilon d_{\chi}(I,I^{\prime}))\mathbb{P}\left(\mathcal{M}(I^{\prime})\in O\right)$
where $d_{\chi}$ is a distance function defined on $\mathcal{I}$. Recent works (Feyisetan et al., 2020; Xu et al., 2020; Li et al., 2023d; Qu et al., 2021) leverage $d_{\chi}$-privacy to safeguard user inputs during both inference and fine-tuning phases. However, as noted by Mattern et al. (2022b), all of the aforementioned DP variants suffer from the linear growth of the privacy budget. A larger privacy budget indicates weaker privacy protection, meaning that as the input length increases, the effectiveness of privacy protection diminishes.
@@ 3 Threat Model
We consider a threat model involving two parties: a server that holds the LLM weights and a client holds user inputs containing privacy spans. Privacy span is defined as a contiguous sequence of tokens within a text that contain private information. The server provides services through an API, enabling the client to transmit inputs and receive responses while maintaining the confidentiality of the LLM weights. The server may be vulnerable to attacks by adversaries seeking to steal privacy information in user inputs. Our task is to protect these private spans in user inputs from being intercepted to the adversaries, even when the adversaries can attack the server directly.
@@ 4 Methodology
To protect privacy, PrivacyRestore transmit the input with privacy spans removed instead of the entire input text to the server. The information in the privacy spans is encrypted as a vector, which is then injected with noise and is also sent to the server. We propose to use activation steering (Li et al., 2023c; Turner et al., 2023; Hernandez et al., 2023) to restore privacy information. Activation steering methods modify the activations of a language model at inference time to predictably alter its behavior. Activation steering is widely used to truthfulness enhancement(Li et al., 2023c), LLMs detoxifying (Li et al., 2024), and sentiment modification (Turner et al., 2023). To our best knowledge, it is first attempt to use activation steering for privacy information restoration. We include the preliminaries of our methodology in Appendix C.
PrivacyRestore operates in two stages, i.e., the preparation stage and the inference stage.
(1) Preparation stage: This stage takes place on the server. We first identify the edited attention heads and train the restoration vectors for each privacy spans. After training, these vectors are released to the clients. The preparation stage is conducted offline, prior to the server beginning to offer its services.
(2) Inference stage: This stage involves collaboration between the client and server. The client constructs a meta vector which is later transmitted to the server along with the input where the privacy spans are removed. The server then performs inference on the input with privacy spans removed and restores those privacy spans via modifying activations.
An overview of PrivacyRestore is shown in Figure 1. Detailed descriptions of the preparation stage and the inference stage are provided in §4.1 and §4.2 respectively. The definitions of all notations used in this paper can be found in Appendix A.
Figure 1: The PrivacyRestore consists of two stages. (1) Preparation Stage. This stage is operated on the server side only and is conducted offline before the server starts offering its services. This stage aims to identify the edited heads and train the restoration vectors. (2) Inference Stage. This stage involves the collaboration between the server and the client. The client need to construct a meta vector by computing a weighted sum of restoration vectors for all privacy spans in the input. A local lightweight model is used to estimate the weight of each privacy span. Then the client transmits the meta vector and the incomplete input with privacy spans removed to the server. Using the meta vector, the server restores the privacy information.
4.1 Preparetion Stage
@@
As pointed by activation steering methods (Li et al., 2023c; Chen et al., 2024), modifying all attention heads in LLMs will degrade overall performance. Inspired by this, we aim to identify the attention heads most relevant to privacy spans.
As shown in upper part of Figure 1, we firstly utilize the probe technique (Alain and Bengio, 2016; Tenney et al., 2019; Belinkov, 2022) to identify the most relevant attention heads for each privacy span. $I_{all}=\{I_{1},...,I_{m}\}$ represents the user inputs in the training set where $m$ is the size of training set. Given a privacy span $s$, $Y_{s}=\{y_{1},...,y_{m}\}$ represents the corresponding labels, where $y_{i}=1$ only if input $I_{i}$ contains privacy span $s$. For each user input $I_{i}$, we record the hidden state of last token on each attention head . We then train a binary classifier for each head, tailored to the privacy span $s$, as the probe. The probe takes the hidden state of last token as input and predicts whether the input contain the privacy span $s$. The probe is formulated as:
$\mathcal{F}^{s}_{h}(\textbf{u}_{h})=\sigma(\theta_{h}^{s}\cdot\textbf{u}_{h}),$ (1)
where $\mathcal{F}^{s}_{h}(\cdot)$ is the probe of privacy token $s$ on head $h$, $\textbf{u}_{h}$ is the hidden state of last token on head $h$, $\theta_{h}^{s}$ is parameters of the probe, and $\sigma(\cdot)$ indicates the sigmoid function. A probe $\mathcal{F}^{s}_{h}(\cdot)$ with higher accuracy indicates a stronger correlation between the head $h$ and the privacy span $s$. Therefore, we select the top $K$ attention heads with highest accuracies for each privacy span.
Subsequently, we introduce a Top-K Heads Selector to identify the common top-K heads set $\mathcal{H}_{c}$ from the individual top-K heads sets of each privacy span. Using different top-K head sets for different privacy spans may suffer the risk of privacy leakage, as an attacker could infer the presence of a specific privacy span based on the characteristics of top-K heads set. Hence, we propose Top-K Heads Selector to combine all different top-K heads sets to construct a common top-K heads set $\mathcal{H}_{c}$ as the edited heads set. To achieve this, we calculate the average score of each head across all privacy spans, selecting the highest $K$ heads to construct the common set. A head receives a positive score if it appears in the top-K head set of a privacy span $s$. The score is related to the accuracy of probe associated with the head. Specifically, if the probe associated with this head yields higher accuracy, the score is higher. By iterating this process across all privacy spans, we can calculate the average score for each head. The detailed algorithm is described in Appendix G.
@@
After identifying the edited heads set, the next step is to train the restoration vectors for each privacy span on the server side. The training objective is to align the predictions given the input with privacy spans removed to be the same as the predictions given an intact input.
For each privacy span $s\in\mathcal{S}$, there is a trainable restoration vector $r_{s}^{h}$ for each head $h$ in the common top-K heads set $\mathcal{H}_{c}$. Restoration vectors of all privacy spans on all heads of $\mathcal{H}_{c}$ form the only trainable parameters $\Theta$ in our method. The LLM weights remain frozen. Our method is plug-and-play and parameter-efficient for training. We fine-tune these restoration vectors using ORPO loss proposed by Hong et al. (2024):
$\quad-\lambda\log\sigma\left(\log\frac{\text{ratio}(a|\hat{I};\Theta)}{\text{ratio}(\hat{a}|\hat{I};\Theta)}\right).$ (3)
where $\hat{I}$ denotes the input with privacy spans removed and $\hat{I}_{all}=\{\hat{I}_{1},\cdots,\hat{I}_{m}\}$ represents the training set of incomplete inputs, $a$ is the initial output give the complete input, $\hat{a}$ is the output given the incomplete input with privacy spans removed and $\lambda$ represents the coefficient. The ORPO loss encourages the model to generate the initial output $a$ rather than the output $\hat{a}$ given the incomplete input. After restoration vectors training, the server will release all restoration vectors to clients.

### v5

@@ Abstract
The widespread usage of online Large Language Models (LLMs) inference services has raised significant privacy concerns about the potential exposure of private information in user inputs. Existing privacy protection methods for LLMs suffer from either insufficient privacy protection with performance degradation, or large inference time overhead. To address these limitations, we propose PrivacyRestore, a plug-and-play method to protect the privacy of user inputs during LLM inference for the client-server scenario. The server first trains restoration vectors for each privacy span type offline and then releases them to the clients. During inference, the client aggregates restoration vectors of all privacy spans in the user query into a meta restoration vector which is later sent to the server to restore information. Before transmission, the client removes all privacy spans in the user query and applies $d_{\chi}$-privacy mechanism to the meta vector for privacy protection. We prove that our method can inherently prevent the linear growth of the privacy budget. We conduct extensive experimental, covering the medical and legal domains, and demonstrate that PrivacyRestore effectively protects private information and maintains acceptable levels of performance and inference efficiency  [^1] We release our code and datasets in https://github.com/wjw136/PrivacyRestore .
@@
Ziqian Zengthanks: Equal contribution.  thanks: Corresponding author  1, Jianwei Wang11footnotemark: 1  1, Junyao Yang11footnotemark: 1  1, Zhengdong Lu1, Haoran Li3, Huiping Zhuang1, Cen Chen1,2 1South China University of Technology 2 Pazhou Laboratory, China 3 Hong Kong University of Science and Technology    zqzeng@scut.edu.cn,    wiwjwilliam@mail.scut.edu.cn
@@ 1 Introduction
Large language models (LLMs) have emerged as powerful tools across various domains (Chen et al., 2023b; Wu et al., ; Wu et al., 2023). However, the widespread use of online LLM inference services has raised significant privacy concerns. When interacting with LLMs deployed on cloud platforms, users’ inputs may contain sensitive data, such as medical records and legal case details. Potential threats may arise when attackers intercept user queries during data transmission, and some advanced adversaries can even hack the cloud service provider. For example, in sensitive domains like medical diagnosis, if a user’s input containing the user’s personal protected health information (PHI), such as “I was previously diagnosed with HIV, and lately I’ve been experiencing fever and diarrhea…” is disclosed to malicious attackers, it may cause privacy concerns.
@@
In this paper, we focus on protecting the private information contained in user inputs during LLM inference. In this setting, the client submits inputs to the server (also known as the service provider) and there is a risk that inputs might be disclosed by attackers. Current methods for protecting user inputs can be mainly divided into two categories: Secure Multi-Party Computation (SMPC) and Differential Privacy (DP). SMPC-based methods (Hao et al., 2022a; Li et al., 2023a; Liang et al., 2024) utilize encryption protocols and algorithms to enable collaborative computation without revealing original data to others. However, SMPC methods have large inference time overheads, making them impractical for real-time applications (Hao et al., 2022b). DP based methods (Feyisetan et al., 2020, 2019; Xu et al., 2020; Bo et al., 2021) apply $d_{\chi}$-privacy (Chatzikokolakis et al., 2013; Alvim et al., 2018) to words and achieve word-level text-to-text privatization. Nevertheless, DP-based methods inevitably degrade the performance of downstream tasks due to noise injection, which is known as the privacy-utility trade-off. Additionally, as the text length grows, word-level privatization will lead to significant performance degradation. This phenomenon is known as the linear growth of the privacy budget in word-level privatization Mattern et al. (2022b). Hence, there is a need to develop privacy-preserving methods which can effectively safeguard the privacy of user inputs while maintaining high-quality outputs, without incurring prohibitive computational costs.
We propose PrivacyRestore, which directly removes privacy spans in user inputs and restores private information via activation steering (Li et al., 2023c; Turner et al., 2023; Hernandez et al., 2023) during model inference. Our method is based on two key phenomena: (a) Users’ private information mostly consists of sensitive attributes and these attributes are commonly confined within specific contiguous token sequences, referred to as “privacy spans”. For eaxmple, in the context of healthcare domain, the private information may commonly refer to symptom descriptions. Consider a medical record which states “I was previously diagnosed with HIV, and lately I’ve been experiencing fever and diarrhea…”, “HIV" and "fever and diarrhea” should be protected as privacy spans. Directly removing these privacy spans (symptom descriptions) can significantly hinder attackers from reconstructing or inferring private information and serves as an effective approach to preventing privacy leakage. (b) Most privacy spans are concentrated in a few majority categories, exhibiting a long-tailed distribution. For instance, in medical diagnosis applications, privacy spans typically relate to symptoms and disease descriptions. Most of the symptoms and disease descriptions appearing in user inputs are concentrated on high-frequency types, such as “fever” and “cold”. We have conducted experiments to demonstrate the long-tailed distribution of privacy spans, as detailed in Appendix K.
@@
PrivacyRestore operates in two stages: the preparation stage and the inference stage. In the preparation stage, the server first identifies the attention heads where the activation steering occurs. Second, each privacy span type is encoded to a vector, known as the restoration vector. This stage is performed entirely offline on the server side. Our method is plug-and-play, requiring only the restoration vectors to be trainable, while keeping the LLM frozen. Once training is complete, these restoration vectors will be released to the client side. In the inference stage, according to the principle of “Information Self-Determination Right”  [^2] https://en.wikipedia.org/wiki/Informational_self-determination (Jasper M C, 2009; Alsenoy et al., 2014) , the users are entitled to identify the privacy spans in their inputs by themselves. After identification, a meta vector is constructed by estimating the importance of each privacy span and calculating a weighted sum of the corresponding restoration vectors. Then, the user submits the incomplete input with the privacy spans removed, along with the meta vector, to the server. The server uses the meta vector to restore the removed privacy spans and generate high-quality outputs.
@@
To prevent the leakage of privacy spans via the meta vector, $d_{\chi}$-privacy mechanisms are applied to the meta vector before transmission at the client side. By applying $d_{\chi}$-privacy to the meta vector instead of words, our method inherently addresses the linear growth issue of privacy budget encountered in word-level privatization (Mattern et al., 2022a). To further prevent privacy leakage through generated outputs, the server should employ sampling-based generation, enabling the output to be protected by the Exponential Mechanism (Utpala et al., 2023a; Mattern et al., 2022c; McSherry and Talwar, 2007). Experimental results demonstrate that our method can effectively protect private information and maintain satisfactory performance and inference efficiency. The contributions are summarized as follows:
@@ 2 Related Works
SMPC-based methods.
Secure Multi-Party Computation (SMPC) uses encryption algorithms to enable secure collaborative computations between the client and server, without revealing the original user inputs to the server.
However, SMPC incurs significant inference time overhead, rendering it impractical for real-time LLM applications (Li et al., 2023a; Liang et al., 2024; Hao et al., 2022a; Liu and Liu, 2023; Zheng et al., 2023b; Gupta et al., 2023; Lu et al., 2024).
DP-based methods.
$d_{\chi}$-privacy mechanisms, a variant of DP, protect user inputs by injecting noise. However, this approach can lead to performance degradation.
Additionally, $d_{\chi}$-privacy becomes less effective as input length increases, due to the linear growth of the privacy budget (Feyisetan et al., 2019; Mattern et al., 2022b; Utpala et al., 2023b; Dwork et al., 2016; Duchi et al., 2013).
Due to the limited space, a detailed introduction of the above works can be found in Appendix B.
@@ 3 Threat Model
We consider a threat model involving two parties: a server that holds the LLM weights and a client holds user inputs containing privacy spans. Privacy span is defined as a contiguous token sequence that contains private information in user inputs, and should be identified by the user itself according to the principle of “Information Self-Determination Right”. The server provides services through an API while maintaining the confidentiality of the LLM weights. Adversaries may intercept privacy spans when users submit their inputs via the API. Relying solely on encryption algorithms is insufficient to prevent privacy leakage, as encrypted inputs will be decrypted on the server and the server itself may be vulnerable. Some advanced adversaries can even hack the server to steal those decrypted user inputs easily (Abrams, 2024; Toulas, 2024). Therefore, our goal is to protect the privacy spans that exist in user inputs from attackers who are capable of stealing user privacy during transmission or even hacking the server to steal user privacy.
@@ 4 Methodology
PrivacyRestore operates in two stages, i.e., the preparation stage and the inference stage, as shown in Figure 1:
(1) Preparation stage: This stage takes place only on the server. Considering the long-tailed distribution of privacy spans, we predefine a core set of privacy span types that covers the majority of them. Next, we identify the edited attention heads required for activation steering during the inference stage. Finally, we train the restoration vector for each privacy span type in the predefined core set. After training, all these vectors are released to the clients. The preparation stage is conducted offline, prior to the server beginning to offer its services.
(2) Inference stage: This stage involves collaboration between the client and server. According to the principle of “Information Self-Determination Right”, the users should identify all privacy spans in their queries by themselves. Then, the client removes all these privacy spans from the queries for privacy protection. For restoration, the client constructs a meta vector according to the removed privacy spans and applies $d_{\chi}$-privacy to the meta vector to prevent privacy leakage. The meta vector, along with the incomplete queries with privacy spans removed, are sent to the server. The server performs inference on the incomplete input and restores information using the meta vector through activation steering.
The preparation and inference stages descriptions are provided in §4.1 and §4.2, respectively. All notation definitions are shown in Appendix A. Backgrounds about the $d_{\chi}$-privacy mechanism and activation steering are shown in Appendix C.
Figure 1: The PrivacyRestore consists of two stages. (1) Preparation Stage. This stage aims to identify the edited heads and train the restoration vectors. We provide a more detailed training set example in Figure 4. (2) Inference Stage. In this stage, the client constructs a meta vector. The server uses the meta vector to restore information during inference on the incomplete query.
4.1 Preparation Stage
@@
As indicated by activation steering methods (Li et al., 2023c; Chen et al., 2024), modifying all attention heads in LLMs will degrade overall performance. Inspired by this, we aim to identify the attention heads most relevant to privacy spans.
As shown in the upper part of Figure 1, we firstly utilize the probe technique (Alain and Bengio, 2016; Tenney et al., 2019; Belinkov, 2022) to identify the most relevant attention heads for each privacy span type. We train a binary classifier for each head, tailored to the privacy span type $c$, as the probe. A probe with higher accuracy indicates a stronger correlation between the head $h$ and the privacy span type $c$. Therefore, we select the top $K$ attention heads with the highest accuracies for each privacy span type $c$ in the predefined core set $\mathcal{C}$.
Using different top-K head sets for different privacy span types may suffer the risk of privacy leakage, as an attacker could infer the presence of a specific privacy span type based on the characteristics of top-K head set. Hence, we propose a Common Top-K Selector to combine all different top-K head sets to construct a common top-K head set $\mathcal{H}_{k}$ as the edited head set. To achieve this, we calculate the average score of each head across all privacy span types in $\mathcal{C}$, selecting the highest $K$ heads to construct the common set. A head receives a positive score if it appears in the top-K head set of a privacy span type $c$. The score is related to the accuracy of the probe associated with the head.
@@
After identifying the edited heads set $\mathcal{H}_{k}$, the next step is to train the restoration vectors for each privacy span type in the predefined core set $\mathcal{C}$.
For each privacy span type $c\in\mathcal{C}$, there is a trainable restoration vector $r_{h}^{c}$ for each head $h$ in the common top-K heads set $\mathcal{H}_{k}$. The restoration vectors constitute the only trainable parameters $\Theta$ in our method, while the LLM weights held by the server remain fixed. Therefore, our method is plug-and-play and parameter-efficient for training.
We fine-tune the restoration vectors using the ORPO loss proposed by Hong et al. (2024), which integrates the supervised fine-tuning process and the preference alignment process. This loss function can guide the model in generating better answers. In our method, we use ORPO loss to train the restoration vectors $\Theta$, ensuring that the outputs generated from inputs without privacy spans and restored using the corresponding restoration vectors, can closely resemble those generated from the intact inputs.
More details of the probe technique, the common top-K head set construction and restoration vectors training process are provided in Appendix D, Appendix E and Appendix F, respectively. After restoration vectors training, the server will release all restoration vectors to clients.